### v3

@@ Abstract
Advancements in clinical treatment are increasingly constrained by the limitations of supervised learning techniques, which depend heavily on large volumes of annotated data. The annotation process is not only costly but also demands substantial time from clinical specialists. Addressing this issue, we introduce the S4MI (Self-Supervision and Semi-Supervision for Medical Imaging) pipeline, a novel approach that leverages the advancements in self-supervised and semi-supervised learning. These techniques engage in auxiliary tasks that do not require labeling, thus simplifying the scaling of machine supervision compared to fully-supervised methods. Our study benchmarks these techniques on three distinct medical imaging datasets to evaluate their effectiveness in classification and segmentation tasks. Notably, we observed that self-supervised learning significantly surpassed the performance of supervised methods in the classification of all evaluated datasets. Remarkably, the semi-supervised approach demonstrated superior outcomes in segmentation, outperforming fully-supervised methods while using 50% fewer labels across all datasets. In line with our commitment to contributing to the scientific community, we have made the S4MI code openly accessible, allowing for broader application and further development of these methods. The code can be accessed at link.
@@ 1 Introduction
Medical imaging analysis is central to clinical assistance with respect to diagnosis, treatment, and planning. With the introduction of deep learning methods, neural networks can now effectively and efficiently support imaging analysis. The automation of this pipeline is becoming of great value for decision-making with respect to patient treatments. However, existing deep-learning techniques rely on the availability of well-structured and annotated medical image data. Unlike natural imaging, annotations are difficult to obtain in the medical domain due to various privacy, cost, and time-related issues [1]. Additionally, the annotation process cannot be crowd-sourced, as annotations in medical imaging require trained clinicians and medical experts to annotate. Consequently, medical imaging datasets are comparatively smaller than natural imaging datasets. Training neural networks from scratch on such a limited amount of data yields sub-par performance. Hence, transfer learning is employed to transfer the knowledge learned in solving source domain tasks to solving target domain tasks. Transfer learning involves initializing the network with weights learned for source domain tasks and fine-tuning target domain data. Transfer learning works best when the source, target dataset image, and output classes are similar [2]; however, this is not always possible in medical imaging.
In recent times, the field of computer vision has witnessed remarkable strides as a result of achievements of learning algorithms such as Deep Metric Learning [3], self-distillation [4], masked image modeling [5], etc. These innovative approaches have made learning meaningful visual representations from unannotated image data possible. Finetuning models from these learners has shown competitive performance [4] gains. Consequently, a growing interest exists in exploring and studying the application of such learning algorithms in medical imaging modalities with limited annotation. These learning algorithms are typically deployed within the framework of self-supervised, semi-supervised, or unsupervised settings. They often involve a two-stage process: 1) a pre-training stage where representations are acquired from unannotated data and 2) a fine-tuning stage using annotated data where the model is refined to align with the specific target task. In a self-supervised setting, pre-training involves an auxiliary task performed without using ground truth labels to learn strong representations. In the case of a semi-supervised setting, the pre-training and finetuning stages can be performed simultaneously. In self-supervised approaches, we focus on DINO (Distillation with NO labels)[4] and CASS (Cross-Architerual Self-supervision) [6]. These techniques leverage joint embedding [7] based architecture to learn useful representations. These approaches essentially learn to map the two views of an image to the same embedding. DINO [4] relies on augmentation to generate two views of the same image, whereas CASS [6] relies on two different neural network architectures for the same image. For the semi-supervised approach, we focus on the cross-architectural, cross-teaching approach [8]. Shifting reliance from the labels (human annotation) to some fully self-supervisory signal or, in part, (semi-supervised) signal could also be understood as a shift to machine supervision from human supervision. Our study shows that machine supervision improves performance with modest fractions of labeled data; the use of these approaches early in the research could save triage time. Machine supervision methods are also more scalable than supervised methods as they are label-independent.
@@
In this work, we compare the performance of state-of-the-art machine supervision approaches against transfer learning on segmentation and classification, two critical medical image analysis tasks, on three challenging datasets of different medical modalities - Histopathology Slide images and skin lesions images. We will also study the effectiveness of machine supervision by varying the amount of annotated data used for fine-tuning. Machine supervision approaches studied are [4, 6] for the purpose of classification, and [8, 9] segmentation.
@@ 2 Methods
2.1 Datasets
To compare transfer learning from ImageNet with machine supervision, we select three datasets representative of the datasets clinicians have to deal with in real life, with varying levels of imbalance and sample sizes.
  • Dermatomyositis Dataset: We have 198 RGB samples from 7 patients, each of 352 by 469 size. With this dataset, we perform multilabel classification. Our task here is to classify cells based on their protein staining into TFH-1, TFH-217, TFH-Like, B cells, and cells that don’t conform to previous cell types as “others” classes. We use the F1 score as our evaluation metric on the test set.
  • Dermofit Dataset: This dataset contains 1300 image samples taken using an SLR camera with the following ten classes: Actinic Keratosis (AK), Basal Cell Carcinoma (BCC), Melanocytic Nevus / Mole (ML), Squamous Cell Carcinoma (SCC), Seborrhoeic Keratosis (SK), Intraepithelial carcinoma (IEC), Pyogenic Granuloma (PYO), Haemangioma (VASC), Dermatofibroma (DF) and Melanoma (MEL). No two images in this dataset are the same size, varying from 205×205 to 1020×1020.
  • ISIC-2017 Dataset: For this dataset, we have 2000 JPEG images classified into three classes - Melanoma, Seborrheic keratosis, and Benign nevi. We use the Recall score as our evaluation metric on the test set.
@@ 2.2.1 Classification pipeline additional details
For the ImageNet supervised classification approach, we used images and their labels as our inputs. For training with xlabels, we only used xof the entire training images and labels while keeping the number of labels per image the same. For the self-supervised approaches DINO and CASS, for pertaining, we start with only images. During the fine-tuning process, we initialize the networks with their pre-trained weights and use corresponding image-label pairs. Similar to the supervised approach, we also fine-tune these architectures for two label fractions, 10% and 100%. Further details can be inferred from our open-sourced code base.
@@ 2.3.3 Approach specific additional details
The semi-supervised model uses a batch size of 16, and each batch consists of eight labeled images and eight unlabeled images. In the last column of Figure 5, we compare the performances between semi- and fully-supervised models in 100% labeled-ratio scenario; for this, we re-adjusted the batch to be 15 labeled images and one unlabeled image to approximate the fully-supervised setting while still retaining the “cross teaching” component so that the loss remains consistent and hence comparable with other semi-supervised label ratios. Additionally, when comparing the performance between fully and semi-supervised models, we adopt the same practice from Luo et al. [8] and use Swin Transformer to compare DEDL with Resnet34 backbone as they have a similar number of parameters
@@
We adopted [10]’s approach for the supervised part of our comparison. For this, we image and mask pairs during training, validation, and testing with a ResNet-34-based U-net.
@@ 3.2 Segmentation
Figure 3: In this figure we present the semi-supervised pipeline as described in 3.2.1. Similar to the classification experiments, we evaluate the segmentation pipeline on three challenging medical image segmentation datasets - the Dermatomyositis, ISIC-2017, and the Dermofit dataset. We use one dataset at a time to train the semi-supervised architecture. Unlike the classification pipeline, semi-supervised learning involves simultaneous learning from labeled and unlabeled data. In Part (b) of Fig. 3, we start by training the data in an unlabeled fashion and during the same iteration labeled data is also passed to the architecture as shown in part (c) of the figure. Predictions from passing inputs of the labeled images yield learned predictions as shown in Part (e). Unsupervised loss is then calculated by comparing the outputs of the CNN and the Transformer (as shown in Part(d) of this figure) using the $\mathcal{L}_{\text{Unsupervised}}$ in Section 3.2.1. This unsupervised loss is then added to the supervised loss denoted by $\mathcal{L}_{\text{Supervised}}$in Section 3.2.1. The supervised loss is calculated against the ground truth as shown in part (e) and (f) in this figure.
@@
Image Segmentation is another important task in medical image analysis. The area of interest – usually containing the pathology – is segmented from the rest of the slide image. We study the performance of transfer learning, semi-supervised, and unsupervised approaches for segmentation. In computer vision, it is widely acknowledged that image segmentation presents a more intricate challenge than image classification. This distinction arises from the fact that image segmentation necessitates the meticulous classification of individual pixels, whereas image classification solely requires making predictions at the image level. Moreover, due to the aforementioned rationale, annotating images for segmentation exhibits a significantly higher complexity level. Therefore, the present study aims to investigate the effectiveness of unsupervised, semi-supervised, and fully supervised techniques across four distinct label fractions instead of the conventional two in classification. In our case, xlabel fractions indicate that we only use xlabels from the dataset. We include the unsupervised approach [9] to study the performance for 0% labels scenario. We compare the performance of DEDL [16], a transfer learning-based approach, against [8], a semi-supervised approach using the Swin transformer-based U-Net model (Figure 3). We choose [16] as it has shown great promise in histopathology imaging and [8], a state-of-the-art approach in leveraging unannotated data for MR segmentation. To ensure a fair comparison, we assess architectures trained with a similar amount of parameters over three datasets and four-label fractions. From [16], we use their best-reported, ResNet-34-based U-Net model. We use thr three data sets as described in Section 2.1. To ensure our results are statistically significant, we conduct all experiments with five different seed values and report the mean values in the 95% confidence interval (C.I) over the five runs. For comparison, we use IoU (Intersection over Union) in line with previous works in the field [10]. IoU or Jaccard index for two images U and V is defined as $IoU(U,V)=\frac{|U\cap V|}{|U\cup V|}$. We present these results in Figure 5.
@@ 3.2.1 Semi-Supervised Approach
1) The labeled data L is evaluated using the average of dice and cross-entropy loss
@@
2) For unlabeled data U cross-teaching strategy is used to cross-supervised between CNN and Transformer: to update U-Net, output from Transformer, $\text{pred}_{\text{Transformer}}$ (vice versa) is used as ”sudo ground truth”:
@@ 3.2.3 Comparision
Figure 5: In this figure, we compare Full, Semi, and Unsupervised architectures for segmentation performance over three datasets. Here, Rn34 represents ResNet-34, while ST represents Swin-U-Net. Since [9] is an unsupervised approach, it does not require any labels for fine-tuning. Hence, we present results for PiCIE using ResNet-34 backbone under 0% label fractions. We observe that the semi-supervised approach outperforms the fully-supervised approach by using 50% fewer per image labels on all three datasets.
@@ 4 Discussion
We observed that the performance of the trained classifiers increased with the introduction of pretraining self-supervised objectives. Comparing the two training regimes, CASS more consistently improved performance over DINO in almost all cases under study. Studying the pixel attribution of the trained methods to understand their decision-making, we observed that CASS usually attributes the relevant pixels more than DINO. Hence, using self-supervision, i.e., additional training even with reduced human supervision, improved performance.
We studied three training regimes for segmentation: supervised, semi-supervised, and unsupervised. While applying the unsupervised approach requires some adaptation and hyperparameter tuning, the dermofit data set consistently shows that semi-supervised outperforms fully-supervised, regardless of labeling percentage. However, we consistently improve performance using a semi-supervised approach – wherein we used labeled and unlabeled data per batch to help the model learn better representation. The result on the dermatomyositis dataset using no labels is close as compared to the supervised and semi-supervised with 10% labels. Between 50% and 70% labeling, the performance of semi-supervised and fully-supervised merges. However, the result on the ISIC-2017 data set with 10% labels indicates fully-supervised performs better than the semi-supervised; however as the percentage of labels increases above 10% semi-supervised outperforms fully-supervised. Additionally, unlike the unsupervised approach, the supervised and the semi-supervised approaches don’t treat this problem as pixel clustering and are less prone to overfitting to the dominant pixel distribution. Since classification requires image-level identification of classes instead of pixel-level identification, segmentation is a more difficult objective. Yet, introducing reduced human supervision with pseudo-labels in the semi-supervised approach improved performance over the supervised and the unsupervised paradigms to the where the semi-supervised approach was able to outperform the fully supervised approach with 50% fewer labels.

### v4

@@ Abstract
Advancements in clinical treatment are increasingly constrained by the limitations of supervised learning techniques, which depend heavily on large volumes of annotated data. The annotation process is not only costly but also demands substantial time from clinical specialists. Addressing this issue, we introduce the S4MI (Self-Supervision and Semi-Supervision for Medical Imaging) pipeline, a novel approach that leverages advancements in self-supervised and semi-supervised learning. These techniques engage in auxiliary tasks that do not require labeling, thus simplifying the scaling of machine supervision compared to fully-supervised methods. Our study benchmarks these techniques on three distinct medical imaging datasets to evaluate their effectiveness in classification and segmentation tasks. Notably, we observed that self-supervised learning significantly surpassed the performance of supervised methods in the classification of all evaluated datasets. Remarkably, the semi-supervised approach demonstrated superior outcomes in segmentation, outperforming fully-supervised methods while using 50% fewer labels across all datasets. In line with our commitment to contributing to the scientific community, we have made the S4MI code openly accessible, allowing for broader application and further development of these methods. The code can be accessed at link.
@@ 1 Introduction
Medical imaging analysis plays a pivotal role in clinical decision-making, aiding in diagnosis, treatment planning, and monitoring. The advent of deep learning has significantly enhanced the capability to analyze medical images both effectively and efficiently, promising to automate aspects of the diagnostic process and thereby augment clinical decision-making. However, the efficacy of these deep learning techniques is heavily contingent upon the availability of well-annotated medical image data. Unlike in natural image processing, obtaining annotations in the medical domain is fraught with challenges due to privacy concerns, high costs, and the extensive time required for expert clinicians to produce accurate annotations [1]. Moreover, the unique requirements for expert knowledge in medical imaging mean that traditional crowd-sourcing for annotations is not viable. This results in medical imaging datasets being significantly smaller than their natural imaging counterparts, leading to suboptimal performance when neural networks are trained from scratch on such limited data volumes. Consequently, transfer learning has become a key strategy, leveraging knowledge acquired from source domain tasks to enhance performance on target domain tasks in medical imaging. Transfer learning typically involves initializing the network with weights from a pre-trained model and subsequently fine-tuning it with target domain data. While transfer learning has shown promise, especially when source and target datasets, as well as their respective output classes, are similar [2], the distinct nature of medical imaging often limits its applicability.
The core motivation of our study is to address these challenges by comparing machine supervision with traditional fully-supervised approaches in the realm of medical imaging. The acquisition of labels in medical imaging is notably costly and time-intensive, a hurdle that machine supervision approaches, with their scalable and label-free nature, aim to overcome. By enabling the development of priors that can outperform supervised methods, machine supervision tackles a critical bottleneck in advancing clinical treatments—the heavy reliance on supervised learning techniques that necessitate extensive annotated data. This study aims to elucidate the efficiencies that machine supervision can introduce, particularly in light of the challenges posed by the need for domain-specific annotation, the limited size of medical datasets due to stringent privacy regulations and high annotation costs, and the indispensable requirement for expert knowledge in annotating medical images.
In recent times, the field of computer vision has witnessed remarkable strides due to achievements in learning algorithms such as Deep Metric Learning [3], self-distillation [4], masked image modeling [5], etc. These innovative approaches have enabled the learning of meaningful visual representations from unannotated image data, with fine-tuning models from these learners showing competitive performance gains [4]. There is a growing interest in exploring and applying these learning algorithms to medical imaging modalities, where annotations are notably scarce. These algorithms, typically deployed in self-supervised, semi-supervised, or unsupervised settings, involve a two-stage process: 1) pre-training to acquire representations from unannotated data, and 2) fine-tuning using annotated data to refine the model for the specific target task.
In the realm of self-supervised learning, our work specifically focuses on DINO (Distillation with NO labels) [4] and CASS (Cross-Architectural Self-supervision) [6] techniques, which are at the forefront of current research. These methods leverage joint embedding [7] based architecture, a cutting-edge approach that facilitates learning meaningful visual representations by aligning two different views of an image to the same embedding space. This strategy is particularly adept at extracting robust features from unannotated data, making it highly relevant to medical imaging analysis. Additionally, for semi-supervised learning, we delve into the cross-architectural, cross-teaching method [8], which represents a significant advancement towards utilizing machine supervision over traditional human annotation. The adoption of such state-of-the-art self/semi-supervised techniques marks a pivotal shift in medical imaging, emphasizing machine supervision’s scalable and label-free advantages. This is crucial in the context of medical imaging, where the procurement of labels is not only cost-prohibitive but also requires extensive time and expertise.
Building upon these considerations, this work introduces the S4MI (Self-Supervision and Semi-Supervision for Medical Imaging) pipeline. This novel semi-/self-supervised learning approach is designed to directly address the aforementioned challenges by minimizing the dependency on extensive labeling efforts. By integrating cutting-edge self-supervised learning algorithms with the latest semi-supervised learning techniques, S4MI aims to significantly improve performance beyond the capabilities of traditional supervised approaches. This initiative represents a significant step towards the development of scalable healthcare solutions, leveraging the untapped potential of machine supervision to achieve and potentially exceed the efficacy of human-supervised methods in medical imaging analysis.
@@
In this work, we evaluate the performance of state-of-the-art machine supervision approaches against traditional transfer learning across two pivotal tasks in medical image analysis: segmentation and classification. Our comparison spans three challenging datasets from distinct medical modalities, including histopathology slide images and skin lesion images. Additionally, we explore the effectiveness of machine supervision by varying the amount of annotated data used for fine-tuning. The machine supervision approaches examined in this study include [4, 6] for classification and [8, 9] for segmentation.
2 Data
To compare transfer learning from ImageNet with machine supervision and show its general applicability, we selected three datasets representative of those clinicians encounter in real life, featuring varying levels of class imbalance and sample sizes.
  • Dermatomyositis: This dataset comprises 198 RGB samples from 7 patients, each image measuring 352 by 469 pixels. With this dataset, we conduct multilabel classification, aiming to classify cells based on their protein staining into TFH-1, TFH-217, TFH-Like, B cells, and cells that do not conform to the aforementioned categories, labeled as ’others.’ We utilize the F1 score as our evaluation metric on the test set.
  • Dermofit: This dataset consists of 1,300 image samples captured with an SLR camera, spanning the following ten classes: Actinic Keratosis (AK), Basal Cell Carcinoma (BCC), Melanocytic Nevus / Mole (ML), Squamous Cell Carcinoma (SCC), Seborrhoeic Keratosis (SK), Intraepithelial Carcinoma (IEC), Pyogenic Granuloma (PYO), Haemangioma (VASC), Dermatofibroma (DF), and Melanoma (MEL). Each image in this dataset is uniquely sized, with dimensions ranging from 205×205 to 1020×1020 pixels.
  • ISIC-2017: Within this dataset, 2,000 JPEG images are distributed among three classes: Melanoma, Seborrheic Keratosis, and Benign Nevi. The evaluation of the test set is based on the Recall score.
@@ 3.1.1 Classification pipeline additional details
For the ImageNet supervised classification approach, we used images and their corresponding labels as our inputs. For training with xlabels, we only used xof the entire training images and labels while keeping the number of labels per image the same. For the self-supervised approaches DINO and CASS, for pertaining, we start with only images. During the fine-tuning process, we initialize the networks with their pre-trained weights and use corresponding image-label pairs. Similar to the supervised approach, we also fine-tune these architectures for two label fractions, 10% and 100%. Further details can be inferred from our open-sourced code base.
@@ 3.2.1 Approach specific additional details
The semi-supervised model uses a batch size of 16, and each batch consists of eight labeled images and eight unlabeled images. In the last column of Figure 5, we compare the performances between semi- and fully-supervised models in 100% labeled-ratio scenario; for this, we re-adjusted the batch to be 15 labeled images and one unlabeled image to approximate the fully-supervised setting while still retaining the “cross teaching” component so that the loss remains consistent and hence comparable with other semi-supervised label ratios. Additionally, when comparing the performance between fully and semi-supervised models, we adopt the same practice from Luo et al. [8] and use Swin Transformer to compare DEDL [10] with Resnet34 backbone as they have a similar number of parameters
@@
We adopted Singh and Cirrone [11]’s approach for the supervised part of our comparison. For this, we utilize image and mask pairs during training, validation, and testing with a ResNet-34-based U-Net[12].
3.3 Common Implementations:
3.3.1 Data-Preprocessing
@@ 4.2 Segmentation
Figure 3: In this figure we present the semi-supervised pipeline as described in 4.2.1. Similar to the classification experiments, we evaluate the segmentation pipeline on three challenging medical image segmentation datasets - the Dermatomyositis, ISIC-2017, and the Dermofit dataset. We use one dataset at a time to train the semi-supervised architecture. Unlike the classification pipeline, semi-supervised learning involves simultaneous learning from labeled and unlabeled data. In Part (b) of Fig. 3, we start by training the data in an unlabeled fashion and during the same iteration labeled data is also passed to the architecture as shown in part (c) of the figure. Predictions from passing inputs of the labeled images yield learned predictions as shown in Part (e). Unsupervised loss is then calculated by comparing the outputs of the CNN and the Transformer (as shown in Part(d) of this figure) using the $\mathcal{L}_{\text{Unsupervised}}$ in Section 4.2.1. This unsupervised loss is then added to the supervised loss denoted by $\mathcal{L}_{\text{Supervised}}$in Section 4.2.1. The supervised loss is calculated against the ground truth as shown in part (e) and (f) in this figure.
@@
Image Segmentation is another important task in medical image analysis. The area of interest – usually containing the pathology – is segmented from the rest of the slide image. We study the performance of transfer learning, semi-supervised, and unsupervised approaches for segmentation. In computer vision, it is widely acknowledged that image segmentation presents a more intricate challenge than image classification. This distinction arises from the fact that image segmentation necessitates the meticulous classification of individual pixels, whereas image classification solely requires making predictions at the image level. Moreover, due to the aforementioned rationale, annotating images for segmentation exhibits a significantly higher complexity level. Therefore, the present study aims to investigate the effectiveness of unsupervised, semi-supervised, and fully supervised techniques across four distinct label fractions instead of the conventional two in classification. In our case, xlabel fractions indicate that we only use xlabels from the dataset. We include the unsupervised approach [9] to study the performance for 0% labels scenario. We compare the performance of DEDL (Data-Efficient Deep Learning framework) [10], a transfer learning-based approach, against [8], a semi-supervised approach using the Swin transformer-based U-Net model (Figure 3).
We choose [10] due to its significant promise in histopathology imaging, and [8] for its state-of-the-art approach to leveraging unannotated data in MR segmentation. The semi-supervised approach [8] trains two segmentation architectures - a Swin transformer-based U-Net and a CNN U-Net (as shown in Fig. 3), while DEDL [10] and unsupervised approach [9] only train one architecture.
To ensure a fair comparison, we make sure the number of overall parameters trained is the same in each case over three datasets and four-label fractions.
In their study, Luo et al. [8] compared the performance of Swin-U-Net trained in semi- and fully-supervised fashion for segmentation. Observing that Swin-U-Net exhibits superior performance in a semi-supervised setting, we extend this comparison to include a semi-supervised Swin-U-Net and a similarly parameterized ResNet-34 based U-Net, trained under both unsupervised and fully-supervised approaches. Furthermore, from [10], we use their best-reported ResNet-34-based U-Net model.
The evaluations are based on three datasets, as detailed in Section 2. To ensure our results are statistically significant, we conduct all experiments with five different seed values and report the mean values in the 95% confidence interval (C.I.) over the five runs. For comparison, we use IoU (Intersection over Union) in line with previous works in the field [11]. IoU or Jaccard index for two images U and V is defined as $IoU(U,V)=\frac{|U\cap V|}{|U\cup V|}$. We present these results in Figure 5.
@@ 4.2.1 Semi-Supervised Approach
[…]
@@
[…]
@@ 4.2.3 Comparision
Figure 5: In this figure, we present the results for the Dermatomyositis dataset, ISIC-2017, and the Dermofit dataset in panels (a), (b), and (c), respectively. We compare the segmentation performance of Full, Semi, and Unsupervised architectures across these datasets, considering different percentages of label fraction (x-axis). Performance is evaluated using the Intersection Over Union (IoU), depicted on the y-axis, to compare results across all three datasets. IoU values range from 0 to 1. The blue bar represents the performance of the unsupervised approach, PiCIE [9], which, by definition, does not require any labels for fine-tuning. Consequently, we present results for PiCIE using 0% label fractions. Remarkably, we observe that the semi-supervised approach surpasses the fully-supervised approach by requiring 50% fewer labels per image across all three datasets.
@@ 5 Discussion
We observed a significant improvement in the performance of trained classifiers with the incorporation of pretraining self-supervised objectives, surpassing methods reliant on fully annotated data. When comparing the two training regimes, CASS consistently outperformed DINO in almost all cases under study. Further analysis into the pixel attribution of the trained methods, to understand their decision-making processes, revealed that CASS typically focuses on relevant pixels more accurately than DINO. Therefore, employing self-supervision—additional training with reduced human supervision—enhanced performance.
We explored three training regimes for segmentation: supervised, semi-supervised, and unsupervised. Although implementing the unsupervised approach necessitates some adaptation and hyperparameter tuning, the Dermofit dataset consistently demonstrates that the semi-supervised approach outperforms the fully-supervised approach, irrespective of the labeling percentage. Notably, we achieved consistent performance improvements using a semi-supervised approach—wherein both labeled and unlabeled data per batch were utilized to enable the model to learn better representations. On the Dermatomyositis dataset, the performance using no labels is closely comparable to that of the supervised and semi-supervised approaches with 10% labels. Performance between semi-supervised and fully-supervised approaches converges when labeling ranges between 50% and 70%. However, on the ISIC-2017 dataset with 10% labels, the fully-supervised approach initially shows superior performance compared to the semi-supervised approach; yet, as the percentage of labels increases beyond 10%, semi-supervised begins to outperform fully-supervised.
Additionally, unlike the unsupervised approach, the supervised and the semi-supervised approaches don’t treat this problem as pixel clustering and are less prone to overfitting to the dominant pixel distribution. Since classification requires image-level identification of classes instead of pixel-level identification, segmentation is a more difficult objective. Yet, introducing reduced human supervision with pseudo-labels in the semi-supervised approach improved performance beyond both the supervised and unsupervised paradigms, to the point where the semi-supervised approach for segmentation outperformed fully-supervised methods while requiring 50% fewer labels across all evaluated datasets.